# C-DiffSET: Leveraging Latent Diffusion for SAR-to-EO Image Translation with Confidence-Guided Reliable Object Generation

### v2

@@ 3.3 Training Strategy for the Diffusion Process
$\mathbf{z}_{y}^{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{y}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\mathcal{N}$ is a Gaussian distribution, $\bm{\epsilon}\in\mathbb{R}^{C\times h\times w}$, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ [18] determines the noise magnitude at each timestep $t$.
@@
where $\tau$ is a margin term ensuring non-negativity of the loss, and $\beta=1$. $\hat{\mathbf{c}}_{t}$ effectively acts as an adaptive weighting factor, allowing the model to focus more on well-aligned regions and reduce penalties in uncertain areas. The $\log$ term serves as a regularizer, preventing $\hat{\mathbf{c}}_{t}$ from collapsing to zero. $\mathcal{L}_{\text{C-Diff}}$ allows C-DiffSET to produce EO outputs of high structural accuracy while mitigating artifacts and hallucinations that can arise in temporally inconsistent regions.
@@ 3.4 Inference Stage for EO Image Prediction
Figure 5: Qualitative SAR-to-EO image translation on the SpaceNet6 dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
@@ 4.1 Datasets
We evaluate our C-DiffSET framework on three publicly available SAR-to-EO datasets: QXS-SAROPT [19], SAR2Opt [80], and SpaceNet6 [56]. These datasets vary in satellite platforms, GSD [61], and SAR polarization modes, allowing us to assess the robustness and generalizability of our approach across diverse real-world scenarios.
@@
Figure 6: Qualitative SAR-to-EO image translation results on the SAR2Opt dataset. Top row: GAN-based methods (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN). Bottom row: LDM-based methods (BBDM, DGDM, cBBDM, and C-DiffSET).
@@
4.2 Experiment Details
All experiments were implemented using PyTorch [42] and conducted on a single NVIDIA A6000 GPU. Each model was fine-tuned for 50,000 iterations, with a 100-step warmup period. We employed the AdamW optimizer [36] with an initial learning rate of $3\times 10^{-5}$ and a weight decay of 0.01. A cosine-annealing scheduler [37] was used to progressively reduce the learning rate at each iteration. For the pretrained LDM, we utilized Stable Diffusion v2.1 [47, 62], and the text encoder was frozen as the CLIP-ViT-H/14 [46, 23] text encoder. For the latent features, we set the spatial resolution parameters as $h=H/8$, $w=W/8$, and the channel dimension as $C=4$. The training noise scheduler was based on DDPM [18] with a total of 1,000 steps. For inference, we employed a more efficient DDIM [58] noise scheduler with total 50 inference steps to accelerate the generation process. Since SAR features were incorporated as additional inputs to the U-Net $\psi$, the weights of the $\psi$’s first convolutional layer were initialized by repeating the original weights across the SAR channels. Additionally, the confidence map was initialized to 1 to ensure stable optimization, starting from the standard $\ell_{2}$ loss, and $\tau=\log 2\pi$ [53] was applied. For the QXS-SAROPT dataset, we used the provided original 256$\times$256-sized patches without modification, with a batch size of 64. For the SAR2Opt and SpaceNet6 datasets, we performed random cropping to 512$\times$512-sized patches and set the batch size to 16. Data augmentation techniques, including random horizontal and vertical flips and random rotations (multiples of 90 degrees), were applied during training to improve generalization. To ensure reproducibility, a random seed of 2,025 was fixed for all experiments. Data was split with 80% used for training and 20% for testing across all experiments. We evaluated the performance of our framework using Fréchet Inception Distance (FID) [16], Learned Perceptual Image Patch Similarity (LPIPS) [78], Spatial Correlation Coefficient (SCC) [83], Structural Similarity Index (SSIM) [65], and Peak Signal-to-Noise Ratio (PSNR).
@@ 4.2 Experiment Details
Table 2: Quantitative comparison of image-to-image translation methods and SET methods on the QXS-SAROPT dataset. Red values represent the best results.
  Pretrained LDM $\mathcal{L}_{\text{C-Diff}}$ SAR2Opt Dataset
[…]
98.98 0.385 0.0014 0.255 15.815
78.14 0.349 0.0029 0.280 16.460
77.81 0.346 0.0035 0.286 16.613
Table 3: Ablation study on the SAR2Opt dataset evaluating the impact of pretrained LDM and confidence-guided diffusion loss.
4.3 Experimental Results
For comparative analysis, we used official implementations for general image-to-image translation methods [24, 84, 69, 33, 71]. For SET-specific methods [31, 28] (marked by $\dagger$), where official codes are generally unavailable due to the specialized nature of the field, we re-implemented each method according to their technical descriptions.
Qualitative comparison. As illustrated in Fig. 1, Fig. 5, and Fig. 6, the GAN-based methods generally struggle with stability during training, leading to prominent artifacts in SET. The LDM-based methods, such as BBDM [33] and cBBDM [28], face challenges due to their reliance on direct diffusion from SAR input $\mathbf{X}$ to EO output $\mathbf{Y}$. This setup makes them particularly susceptible to local spatial misalignments, producing blurred and incoherent results. DGDM [71] seeks to improve the initial stage of diffusion through a deterministic approach, employing a lightweight translation network to create a translated EO latent from SAR latent. However, the simplistic nature of this translation network fails to capture the intricate EO features, resulting in suboptimal initial latents for the diffusion process. In contrast, our C-DiffSET leverages pretrained LDM as foundational model, effectively addressing alignment issues via a confidence-guided diffusion loss, leading to yield sharper and more structurally coherent EO images, and to outperform both GAN-based and diffusion-based baselines in visual fidelity. Notably, as shown in Fig. 6, slight discrepancies between our generated EO images and the target EO images can occur. These differences arise from our model’s emphasis on structural alignment with the input SAR image, prioritizing SAR-EO correspondence over exact visual matching with the target EO.
Quantitative evaluation. In the quantitative assessment (see Table 1 and Table 2), the GAN-based methods yield lower PSNR, SSIM, and SCC scores, indicating their propensity for artifacts and limited spatial coherence. The LDM-based methods, BBDM and cBBDM, also perform inadequately due to their direct diffusion setup from SAR to EO, which exacerbates pixel-wise misalignment issues and inflates LPIPS and FID scores. The DGDM incorporates a lightweight translator network to initialize the diffusion process, but its simplistic structure fails to encapsulate detailed EO features, resulting in lower PSNR and SSIM scores. In contrast, our C-DiffSET achieves superior performance across all metrics, enabled by pretrained LDM and the confidence-guided diffusion loss, which jointly tackles SAR-specific noise and alignment challenges. This adaptation allows the C-DiffSET to attain the highest PSNR, SSIM, and SCC values, along with the lowest LPIPS and FID scores, highlighting its enhanced fidelity and perceptual quality in SET.
@@ 4.4 Ablation Studies
Impact of pretrained LDM and $\mathcal{L}_{\text{C-Diff}}$. To assess the contribution of the pretrained LDM and the confidence-guided diffusion loss $\mathcal{L}_{\text{C-Diff}}$, we conducted an ablation study on the SAR2Opt dataset, as shown in Table 3. The pretrained LDM provides strong initialization for EO generation, improving the baseline PSNR and SSIM scores when compared to training from scratch. Furthermore, incorporating $\mathcal{L}_{\text{C-Diff}}$ somewhat enhances these metrics and lowers LPIPS and FID values, improving both perceptual fidelity and structural consistency. Fig. 7 shows confidence maps generated on the training dataset at timestep $t=T/2$ (midpoint of the denoising process), highlighting the areas of temporal discrepancy in SAR-EO pairs. This loss enables the U-Net $\psi$ to down-weight uncertain regions where objects are temporally misaligned across modalities, thereby reducing artifacts and ensuring coherent EO outputs.
@@ A.1 Additional Qualitative Comparisons
Fig. 11, Fig. 12, Fig. 13, Fig. 14, Fig. 15, Fig. 16, and Fig. 17 provide additional qualitative comparisons of SAR-to-EO image translation results on the QXS-SAROPT [50], SAR2Opt [80], and SpaceNet6 [56] dataset. The GAN-based methods, including Pix2Pix [24] and CycleGAN [84], exhibit severe artifacts due to inherent unstable training in GAN frameworks. Although CFCA-SET [31] and StegoGAN [69] mitigate some of these artifacts, they still produce visually inconsistent results, often failing to preserve fine-grained structural details. Among the LDM-based approaches, BBDM [33] and cBBDM [28] generate smoother outputs and struggle with oversimplified textures and lack of structural alignment due to their direct diffusion process. DGDM [71], on the other hand, relies on an initial translator network to generate SAR-to-EO latents; however, the simplicity of this translator leads to poorly initialized latents, resulting in entirely unrealistic outputs. In contrast, our proposed C-DiffSET effectively addresses these limitations, producing visually coherent and structurally accurate EO images that are closely aligned with the target EO images.
A.2 Progressive Denoising Visualization
A.3 Analysis of VAE Reconstruction and Upper Bound Performance
In Fig. 4, we evaluate the reconstruction quality of the VAE encoder $\mathcal{E}_{\text{vae}}$ and decoder $\mathcal{D}_{\text{vae}}$ from the LDM by visualizing their outputs for target EO images and SAR images with varying levels of speckle noise. Our proposed C-DiffSET framework, built upon the LDM architecture, predicts target EO features $\mathbf{z}_{y}=\mathcal{E}_{\text{vae}}(\mathbf{Y})$ embedded in the latent space through the VAE encoder. Thus, the VAE-decoded reconstruction of the target EO feature, $\mathcal{D}_{\text{vae}}(\mathcal{E}_{\text{vae}}(\mathbf{Y}))$, serves as an upper bound for our C-DiffSET, representing the maximum achievable quality given the pretrained LDM. Table 4 quantitatively compares the performance of C-DiffSET with this upper bound. While our framework delivers strong results, especially in terms of perceptual metrics such as LPIPS and FID, the gap to the upper bound highlights the inherent limitations in SAR-to-EO image translation due to noise, misalignments, and the complexities of SAR data. This analysis underscores the potential for future work to narrow this gap, advancing SAR-to-EO image translation toward the limits defined by the upper bound.
  Methods First 50 samples from the SpaceNet6 Dataset
@@ A.3 Analysis of VAE Reconstruction and Upper Bound Performance
$\mathcal{D}_{\text{vae}}(\mathcal{E}_{\text{vae}}(\mathbf{Y}))$ 43.57 0.049 0.1631 0.783 28.633

### v3

@@ 3.3 Training Strategy for the Diffusion Process
$\mathbf{z}_{y}^{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{y}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\mathcal{N}$ is a Gaussian distribution, $\bm{\epsilon}\in\mathbb{R}^{h\times w\times C}$, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ [18] determines the noise magnitude at each timestep $t$.
@@
where $\tau$ is a margin term ensuring non-negativity of the loss. $\hat{\mathbf{c}}_{t}$ effectively acts as an adaptive weighting factor, allowing the model to focus more on well-aligned regions and reduce penalties in uncertain areas. The $\log$ term serves as a regularizer, preventing $\hat{\mathbf{c}}_{t}$ from collapsing to zero. Empirically, we found that setting $\beta=1$ yields the best performance, while $\beta=0$ reduces $\mathcal{L}_{\text{C-Diff}}$ to a standard $\ell_{2}$ (MSE) loss, lacking adaptive weighting. $\mathcal{L}_{\text{C-Diff}}$ allows C-DiffSET to produce EO outputs of high structural accuracy while mitigating artifacts and hallucinations that can arise in temporally inconsistent regions.
@@ 3.4 Inference Stage for EO Image Prediction
Figure 5: Visual comparison of SET results on the SpaceNet6 and SAR2Opt datasets. 1st rows: GAN-based (Pix2pix, CycleGAN, CFCA-SET, and StegoGAN) methods. 2nd rows: LDM-based (BBDM, ControlNet, Uni-ControlNet, DGDM, cBBDM, and C-DiffSET) methods.
@@ 4.1 Datasets
We evaluate our C-DiffSET framework on three publicly available SAR-to-EO datasets: QXS-SAROPT [19], SAR2Opt [83], and SpaceNet6 [56]. These datasets vary in satellite platforms, GSD [61], and SAR polarization modes, allowing us to assess the robustness and generalizability of our approach across diverse real-world scenarios. The SAR images in these publicaly released datasets are provided as magnitude-only representation, containing real-valued intensity without phase information.
@@
4.2 Experiment Details
All experiments were implemented using PyTorch [43] and conducted on a single NVIDIA A6000 GPU. Each model was fine-tuned for 50,000 iterations, with a 100-step warmup period. We employed the AdamW optimizer [37] with an initial learning rate of $3\times 10^{-5}$ and a weight decay of 0.01. A cosine-annealing scheduler [38] was used to progressively reduce the learning rate at each iteration. For the pretrained LDM, we utilized Stable Diffusion v2.1 [47, 63], and the text encoder was frozen as the CLIP-ViT-H/14 [46, 23] text encoder. For the latent features, we set the spatial resolution to $h=H/8$ and $w=W/8$, and the channel dimension to $C=4$. The training noise scheduler was based on DDPM [18] with a total of 1,000 steps. For inference, we employed an efficient DDIM [58] noise scheduler with total 50 inference steps to accelerate the generation process. We evaluated the performance of our framework using Fréchet Inception Distance (FID) [17], Learned Perceptual Image Patch Similarity (LPIPS) [80], Spatial Correlation Coefficient (SCC) [86], Structural Similarity Index (SSIM) [66], and Peak Signal-to-Noise Ratio (PSNR).
4.3 Experimental Results
For comparative analysis, we used official implementations for general image-to-image translation methods [24, 87, 70, 33, 72, 77, 82]. For LDM-based methods, including our C-DiffSET, it should be noted that we ensured fair comparison by initializing all models with the same Stable Diffusion v2.1 weights (VAE, U-Net). For SET-specific methods [31, 28, 73] (marked by $\dagger$), where their official codes are often unavailable due to this specific field, we re-implemented each method according to their technical descriptions.
Qualitative comparison. As shown in Fig. 1 and Fig. 5, the GAN-based methods generally struggle with stability during training, leading to prominent artifacts in SET. The LDM-based methods, such as BBDM [33] and cBBDM [28], face challenges due to their reliance on direct diffusion from SAR input $\mathbf{X}$ to EO output $\mathbf{Y}$. This setup makes them particularly susceptible to local spatial misalignments, producing blurred and incoherent results. DGDM [72] seeks to improve the initial stage of diffusion through a deterministic approach, employing a lightweight translation network to create a translated EO latent from SAR latent. However, the simplistic nature of this translation network fails to capture the intricate EO features, resulting in suboptimal initial latents for the diffusion process. Furthermore, ControlNet-based approaches [77, 82] keep the pretrained U-Net weights frozen and only introduce SAR conditions at the decoder stage, limiting their ability to fully integrate SAR structural information. This constraint results in weaker feature adaptation and reduced robustness to SAR-induced artifacts. In contrast, our C-DiffSET leverages pretrained LDM as foundational model, effectively addressing alignment issues via a confidence-guided diffusion loss, leading to yield sharper and more structurally coherent EO images, and to outperform both GAN-based and diffusion-based baselines in visual fidelity.
Quantitative evaluation. In Table 1 and Table 2, the GAN-based methods benefit from the richer polarization diversity in SpaceNet6 dataset (4-channel full-polarization SAR data), yielding higher SSIM and SCC scores compared to 1-channel single-polarization datasets. However, for the other peceptual quality metrics such as FID and LPIPS, the GAN-based methods yield lower performance across datasets, indicating their inherent limitations in handling the SET task. The LDM-based methods, BBDM and cBBDM, also perform inadequately due to their direct diffusion setup from SAR to EO, which exacerbates the pixel-wise misalignment problem and inflates LPIPS and FID scores. The DGDM incorporates a lightweight translator network to initialize the diffusion process, but its simplistic structure fails to encapsulate detailed EO features, resulting in lower PSNR and SSIM scores. In contrast, our C-DiffSET achieves superior performance across all metrics, enabled by the pretrained LDM and confidence-guided diffusion loss, which jointly tackle SAR-specific noise and alignment challenges. This adaptation allows C-DiffSET to attain the highest PSNR, SSIM, and SCC values, along with the lowest LPIPS and FID scores, highlighting its enhanced fidelity and perceptual quality in SET.
@@ 4.3 Experimental Results
Table 2: Quantitative comparison of image-to-image translation methods and SET methods on the QXS-SAROPT dataset.
@@ 4.4 Ablation Studies
Impact of pretrained LDM and $\mathcal{L}_{\text{C-Diff}}$. We ablate the contribution of the pretrained LDM and the confidence-guided diffusion loss $\mathcal{L}_{\text{C-Diff}}$ for the SAR2Opt and SpaceNet6 datasets, and show the results in Table 3. The pretrained LDM provides strong initialization for EO generation, improving the baseline PSNR and SSIM scores when compared to training from scratch. Even though the LDM-based methods and our C-DiffSET utilize the same pretrained LDM weights, our C-DiffSET outperformed the others by embedding conditioned SAR and target EO images into the shared latent space, leading to stable convergence. Furthermore, $\mathcal{L}_{\text{C-Diff}}$ helps enhancing the SSIM and PSNR metrics and lowering LPIPS and FID values, thus improving both perceptual fidelity and structural consistency. Fig. 6 shows confidence maps generated on the training dataset at timestep $t=T/2$ (midpoint of the denoising process), highlighting the areas of temporal discrepancy in SAR-EO pairs. This loss enables the U-Net $\psi$ to down-weight uncertain regions where objects are temporally misaligned across modalities, thereby reducing artifacts and ensuring coherent EO outputs. Further results and detailed analysis can be found in the Supplemental Material.
@@ 4.5 Limitations
  Pretrained LDM Loss function SAR2Opt / SpacNet6 Dataset
[…]
MSE 98.98 / 60.26 0.39 / 0.23 0.001 / 0.011 0.26 / 0.43 15.82 / 18.16
MSE 78.14 / 40.62 0.36 / 0.16 0.003 / 0.014 0.28 / 0.52 16.46 / 20.29
C-Diff 77.81 / 37.44 0.34 / 0.14 0.004 / 0.015 0.29 / 0.57 16.61 / 21.02
Table 3: Ablation studies on the SAR2Opt and SpaceNet6 dataset evaluating the impact of pretrained LDM and confidence-guided diffusion (C-Diff) loss.
@@ A.1 Additional Qualitative Comparisons
Fig. 16, Fig. 12, Fig. 13, Fig. 14, Fig. 15, Fig. 17, Fig. 18, Fig. 19, and Fig. 20 provide additional qualitative comparisons of SAR-to-EO image translation results on the QXS-SAROPT [50], SAR2Opt [83], and SpaceNet6 [56] datasets. The GAN-based methods, including Pix2Pix [24] and CycleGAN [87], exhibit severe artifacts due to the inherent instability of the training process within the GAN frameworks. Although CFCA-SET [31] and StegoGAN [70] mitigate some of these artifacts, they still produce visually inconsistent results, often failing to preserve fine-grained structural details. Among the LDM-based approaches, BBDM [33] and cBBDM [28] generate smoother outputs and struggle with oversimplified textures and lack of structural alignment due to their direct diffusion process. DGDM [72] relies on an initial translator network to generate SAR-to-EO latents; however, the simplicity of this translator leads to poorly initialized latents, resulting in entirely unrealistic outputs. ControlNet-based approaches [77, 82] exhibit similar limitations, as they keep the pretrained U-Net weights frozen and inject SAR conditions only in the decoder stage, failing to fully propagate the structural information of SAR throughout the denoising process. In contrast, our proposed C-DiffSET effectively addresses these limitations, producing visually coherent and structurally accurate EO images that are closely aligned with the target EO images.
A.2 Analysis of Domain Gap and Upper Bound Performance
In Fig. 4 of the main paper, we evaluate the reconstruction quality of the VAE encoder $\mathcal{E}_{\text{vae}}$ and decoder $\mathcal{D}_{\text{vae}}$ from the LDM by visualizing their outputs for target EO images and SAR images with varying levels of speckle noise.
Embedding into a shared latent space. In Fig. 7, additional VAE reconstruction results demonstrate that SAR and EO images are embedded into the same latent space. Notably, VAE-reconstructed SAR images preserve structural information while reducing speckle noise, enhancing their suitability as conditioning inputs for SET task.
Domain gap analysis. Fig. 7, Table 4, Table 5, and Table 6 show that the VAE, trained on natural image datasets, effectively reconstructs the RGB bands of EO images. This indicates that the domain gap between natural images and EO imagery in the RGB spectrum is relatively small, enabling the pretrained LDM to generalize well to EO data. This observation further supports our approach of leveraging a large-scale pretrained diffusion model for the SET task.
Figure 7: Results of applying the VAE encoder and decoder from LDM to EO and SAR images. These results suggest that the LDM VAE facilitates a shared latent representation between SAR and EO domains, supporting robust cross-domain image translation.
  Methods SpaceNet6 Test Dataset
@@ A.2 Analysis of Domain Gap and Upper Bound Performance
$\mathcal{D}_{\text{vae}}(\mathcal{E}_{\text{vae}}(\mathbf{Y}))$ 17.64 0.047 0.1417 0.800 28.543
@@
  Methods SAR2Opt Test Dataset
[…]
C-DiffSET (Ours) 77.81 0.346 0.0035 0.286 16.613
$\mathcal{D}_{\text{vae}}(\mathcal{E}_{\text{vae}}(\mathbf{Y}))$ 28.99 0.095 0.1583 0.667 25.678
Table 5: Comparison of C-DiffSET performance with the VAE upper bound on the SAR2Opt dataset.
  Methods QXS-SAROPT Test Dataset
[…]
C-DiffSET (Ours) 18.15 0.293 0.0108 0.372 18.077
$\mathcal{D}_{\text{vae}}(\mathcal{E}_{\text{vae}}(\mathbf{Y}))$ 9.18 0.064 0.2597 0.794 29.497
Table 6: Comparison of C-DiffSET performance with the VAE upper bound on the QXS-SAROPT dataset.
  Types Methods Params. (M) FLOPs (G) Memory (MB) Time (s)
GANs Pix2Pix [24] 54.41 24.22 464.12 0.06
CycleGAN [87] 7.84 140.43 398.38 0.08
SAR-SMTNet [73] 2.15 615.40 2626.98 0.22
CFCA-SET [31] 26.80 98.98 431.86 0.10
StegoGAN [70] 13.15 227.49 461.14 0.11
LDMs BBDM [33] 949.56 2122.44 6147.88 3.13
ControlNet [77] 1312.72 2231.03 7567.83 4.50
Uni-ControlNet [82] 1519.18 2295.20 8382.00 4.95
DGDM [72] 959.03 2161.25 6184.45 1.47
cBBDM [28] 949.58 2122.49 6147.93 3.15
C-DiffSET 949.58 2122.49 6148.09 3.27
Table 7: Comparative analysis of C-DiffSET with other methods by parameters, FLOPs, memory usage, and inference time.
Figure 8: Visual comparison of SET results with and without C-Diff loss. The MSE loss corresponds to $\beta=0$, indicating no confidence weighting.
@@
  Pretrained LDM Loss function QXS-SAROPT Dataset
[…]
MSE 29.04 0.407 0.0006 0.279 14.647
MSE 19.99 0.297 0.0094 0.364 17.736
C-Diff 18.15 0.293 0.0108 0.372 18.077
Table 8: Ablation studies on the QXS-SAROPT dataset evaluating the impact of pretrained LDM and confidence-guided diffusion (C-Diff) loss.
  Text Prompt SpaceNet6 Dataset
[…]
“  ” (Null text, $\varnothing$) 79.01 0.351 0.0027 0.273 16.237
“Eletro-Optical Image” 77.81 0.346 0.0035 0.286 16.613
Table 9: Ablation studies on the text prompts.
Upper bound comparison. Our proposed C-DiffSET framework, built on the LDM architecture, predicts the target EO features $\mathbf{z}_{y}=\mathcal{E}_{\text{vae}}(\mathbf{Y})$ embedded in the latent space through the VAE encoder. The VAE-decoded reconstruction of the target EO feature, $\mathcal{D}_{\text{vae}}(\mathbf{z}_{y})=\mathcal{D}_{\text{vae}}(\mathcal{E}_{\text{vae}}(\mathbf{Y}))$, serves as an upper bound for our C-DiffSET, representing the maximum achievable quality given the pretrained LDM. Table 4, Table 5, and Table 6 quantitatively compare the performance of C-DiffSET against this upper bound. Although our framework achieves strong results, particularly in perceptual metrics such as LPIPS and FID, the gap to the upper bound highlights the inherent challenges in SAR-to-EO translation, including noise, misalignment, and the structural complexity of SAR data. This analysis underscores the potential for future research to bridge this gap, pushing SET closer to the theoretical upper limit.
A.3 Computational Complexity
We compare the number of parameters, FLOPs, memory usage, and inference time for 512$\times$512 images in Table 7. As noted, the LDM-based methods are more complex than the GAN-based ones. The time-performance trade-off is in Fig. 9.
A.4 Progressive Denoising Visualization